OpenStudy (emilyngiron):

Please help! I have a test and a teacher that cant teach. Suppose a biologist captured and tagged 15 fish from a lake. The tagged fish were released back into the lake. A sample was taken some time later. In that sample there were 800 fish and only 5 were tagged. Predict how many fish in total are in the lake.

OpenStudy (giitter):

Im sorry we do not answer test questions only homework questions

OpenStudy (giitter):

I truly am sorry

OpenStudy (emilyngiron):

this isnt the actual test. its supposed to prepare me but my teacher doesnt teach us anything. he barely speaks english.

OpenStudy (emilyngiron):

@GIitter

OpenStudy (giitter):

Aww

OpenStudy (giitter):

hang on

OpenStudy (dumbcow):

Use a proportion: \[\frac{sample -total}{sample -tagged} = \frac{total}{total-tagged}\] \[\frac{800}{5} = \frac{x}{15}\] \[x = 15(\frac{800}{5}) = 2400\]

OpenStudy (emilyngiron):

okay that was simple. it was different tan the example on the page